# Global Balance and Systemic Risk in Financial Correlation Networks

### v1

@@ 4.3 Approximating Global Balance and Condition Numbers in Correlation Networks
We now show that the ratio ${\mathscr{R}}({\bf A})$ is more typically less than $1$ for a large set of significant matrices, in particular for the empirical correlation matrices we are interested in. We assume from now on that the matrix ${\bf A}$ is positive semidefinite so that all its eigenvalues $\lambda_{i}$ are nonnegative: $0\leq{\lambda}_{N}\leq{\lambda}_{N-1}\leq\dots\leq{\lambda}_{2}\leq{\lambda}_{1}$. We refer to ${\lambda}_{1}$ as the spectral radius and to ${\lambda}_{1}-{\lambda}_{2}$ as the spectral gap.
We first express ${\mathscr{R}}({\bf A})$ as a function of the intervals between pairs of eigenvalues, as shown in the following.
Proposition 5.
The ratio ${\mathscr{R}}({\bf A})$ between the condition numbers of the matrices $e^{-{\bf A}}$ and $e^{-|{\bf A}|}$, in the Schatten $p-$norm with $p=1$, can be expressed as
@@
where ${\lambda}_{N}\leq{\lambda}_{N-1}\leq\dots\leq{\lambda}_{1}$ and $\overline{\lambda}_{N}\leq\overline{\lambda}_{N-1}\leq\dots\leq\overline{\lambda}_{1}$ are the eigenvalues of $\bf A$ and $\bf|A|$, and $\cosh(\cdot)$ denotes the hyperbolic cosin function.
Proof.
The statement immediately follows from $\left(\sum_{i=1}^{N}e^{-\lambda_{i}}\right)\left(\sum_{i=1}^{N}e^{\lambda_{i}}\right)=N+2\sum_{j<i}\cosh(\lambda_{i}-\lambda_{j})$. ∎
Proposition 5 emphasizes the dependence of ${\mathscr{R}}({\bf A})$ on the intervals $\Delta\lambda_{ij}=\lambda_{i}-\lambda_{j}$ between pairs of eigenvalues. It highlights the relevant role of the spectral gap in large correlation networks based on the correlation matrix ${\bf C}$.
@@
${\mathscr{R}}({\bf A})\approx\frac{N+2\sum_{j=2}^{N}\cosh\Delta\lambda_{1j}}{N+2\sum_{j=2}^{N}\cosh\Delta\overline{\lambda}_{1j}}\approx\frac{N+2(N-1)\cosh\lambda_{1}}{N+2(N-1)\cosh\overline{\lambda}_{1}}\approx\frac{e^{\lambda_{1}}}{e^{\overline{\lambda}_{1}}}$ (20)
In general, for any signed network, the inequality $\lambda_{1}\leq\overline{\lambda}_{1}$ holds. The spectral radius of the matrix ${\bf A}$ is strictly smaller than that of the matrix $|{\bf A}|$ if and only if the signed network $G$ is strictly unbalanced, that is neither balanced nor antibalanced, i.e. ${\lambda}_{1}<\overline{\lambda}_{1}$ if and only if $G$ is strictly unbalanced (see [46], Theorem 3.9, and [50]). This implies that the approximated value in Eq. (20) is less than $1$ for strictly unbalanced networks. Let us observe that, since $\sum_{j=2}^{N}\lambda_{j}=N-\lambda_{1}$ and $\sum_{j=2}^{N}\overline{\lambda}_{j}=N-\overline{\lambda}_{1}$, we have $\sum_{j=2}^{N}\Delta\lambda_{1j}=N(\lambda_{1}-1)<N(\overline{\lambda}_{1}-1)=\sum_{j=2}^{N}\Delta\overline{\lambda}_{1j}$, but this does not guarantees that $\Delta\lambda_{1j}<\Delta\overline{\lambda}_{1j},\ \forall j$. However, if ${\lambda}_{1}\gg{\lambda}_{j}$ and $\overline{\lambda}_{1}\gg\overline{\lambda}_{j}$, for $j=2,\dots,N$, we can also assume $\Delta{\lambda}_{1j}\approx{\lambda}_{1}$ and $\Delta\overline{\lambda}_{1j}\approx\overline{\lambda}_{1}$. This justifies the second approximation in Eq. (20). As a further consequence of the same equation, we can conclude that for correlation networks with large spectral gap the value of the ratio ${\mathscr{R}}({\bf A})$ can be approximated with the value of the global balance $\kappa(G)$, that is ${\mathscr{R}}({\bf A})\approx\kappa(G)$. This effectiveness of the approximation is confirmed by the results in the empirical analysis.
@@
Our interpretation of this result is that the presence of negative edges, that is negative correlations, mitigates the propagation of a perturbation shock within the network. Therefore, an unbalanced graph can mitigate and absorb the perturbations from one node better than a balanced graph. In a strictly unbalanced graph, we expect that the oscillations produced by a single node cannot be amplified to a large extent due to the lower condition number of the adjacency matrix. Conversely, in a balanced graph, where there are two disjoint communities and all closed walks are positive, a perturbation originating from one node is amplified travelling through the network. This result shows that a balanced graph is more unstable and sensitive than an unbalanced one and, at the same time, it is more reliable in transmitting information. This supports the interpretation of network global balance in terms of reliability in transmitting information, as discussed in subsection 4.1. We conclude this section with a numerical example.
@@
$A=\left[\begin{array}[]{cccc}1&0.76442737&0.33531137&0.2356325\\ 0.76442737&1&0.05137479&-0.3490933\\ 0.33531137&0.05137479&1&0.7348348\\ 0.2356325&-0.3490933&0.7348348&1\\ \end{array}\right].$ (21)
The network associated with this correlation matrix is a weighted version of the graph $K_{4}$ (d) in Fig. 4.1, third line. It is a strictly unbalanced network and its global balance is $\kappa(G)=0.9616991$. The local balances of the four nodes are collected in the vector $[0.9596102,0.9519667,0.9832867,0.9526123]$.
Let us first note that by increasing the initial state ${\bf x}_{0}=[1,1,1,1]^{T}$, for example, by the error $\varepsilon{\bf e}_{1}$ with $\varepsilon=2$, the increase in the asymptotic state of node $1$, that is ${\bf x}_{\infty,1}^{\star}-{\bf x}_{\infty,1}=\varepsilon[e^{\bf A}]_{11}$, is equal to $7.640996$ for the signed network and to $7.962604$ for the unsigned network. The difference for a unitary error, equal to $[e^{|{\bf A}|}]_{11}-[e^{\bf A}]_{11}=0.160804$, increases linearly with $\varepsilon$ and the ratio $[e^{\bf A}]_{11}/[e^{|{\bf A}|}]_{11}=0.9596102$ is exactly the local balance of node $1$. The eigenvalues of ${\bf A}$ are $\lambda_{1}=1.95627077$ $\lambda_{2}=1.70464321$. $\lambda_{3}=0.31667046$ and $\lambda_{4}=0.02241557$, while the eigenvalues of $|{\bf A}|$ are $\overline{\lambda}_{1}=2.23949471$ $\overline{\lambda}_{2}=1.27400244$. $\overline{\lambda}_{3}=0.43924268$ and $\overline{\lambda}_{4}=0.04726018$. The two exact condition numbers are $\mathscr{K}(e^{-{\bf A}})=30.37852$ and $\mathscr{K}(e^{-{{|\bf A}|}})=30.88739$, and their exact ratio is therefore ${\mathscr{R}}({\bf A})=0.9835251$. For such a low-dimensional correlation matrix, the difference between the two condition numbers is small. Nevertheless it evidences the fact that the balanced network shows a greater dependence on the initial data than the unbalanced network.
This numerical example can be extended to a statistical analysis of the relationship between the global balance $\kappa(G)$ of the network $G$ and the ratio of the two condition numbers ${\mathscr{R}}({\bf A})$ in the case of randomly generated correlation matrices. We refer to the A for further details.
@@ 4.4 Interpretation in terms of systemic risk measure
Eq. (22) shows that $\lambda_{i}$ represents the portion of the total variance associated with the $i-{th}$ principal portfolio ${\bf\phi}_{i}$, derived from PCA analysis, as defined in [51]. Each component $\lambda_{i}{\bf\phi}_{i}{\bf\phi}_{i}^{T}$ represents a new portfolio out of the original stocks, which is orthogonal with the others. Then the corresponding eigenvalue measures its contribution to the total risk.
@@
where $1\leq H\leq N$ and $p\in{\mathbb{N}}$. Note that the AMRI reduces to the condition number in norm $2$, $\lambda_{1}/\lambda_{N}$, for $H=1$. This measure has been shown to belong to the broad class of so-called proper measures of connectedness ([53]) and have been used to empirically analyze financial datasets in [54]. [^3] Let us observe that, in graphs theory, the quantity $\frac{1}{N}{\rm Tr}\,\bf{C}^{p}=\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}^{p}$ represents the $p-$moment of the adjacency matrix $\bf C$ and, in a positive unweighted network, represents the mean of the number of $p-$cycles around the nodes, that is the average number of $p-$cycles to which a node in the network belongs.
The local and global balance indices introduced in Eqs. (9) and (10) are defined in a way completely independent of both the class of systemic risk measures based on the larger eigenvalues and the class based on the smaller eigenvalues, since they imply a comparison between the correlation matrix and its entry-wise absolute value.
However, the interpretation proposed in the previous sections suggests that a bridge can be established between the domain of balance in signed networks and the domain of systemic risk measures for asset correlation networks. Indeed, the relationship of the balance index with the condition numbers of the diffusion problem and with the ratio ${\mathscr{R}}({\bf A})$ defined in Eq. (16) indicates that they are candidates to be indicators of systemic risk as a relative measure of the effect of correlation signs on the spread of information in the underlying network. [^4] It is also possible to show that the global balance trivially satisfies the four properties required in [53] for a proper measure of connectedness. As a consequence, they can be used to detect the greater or lesser stability/sensitivity of the portfolio behavior with respect to the maximum stability/sensitivity associated with the balanced one, which typically has a higher condition number and, consequently, are less diversified and more exposed to systemic risk. At this point, we now want to explore the ability of this new indicator to reveal systemic events on empirical correlation networks and compare its capabilities with that of other known measures in the literature.
@@ 5 Empirical analysis
We test the results of Sections 3 and 4 on two different databases. We focus the analysis on the weighted version of the correlation-based signed network. However, we also compare the results with the binary network. The binarization is obtained by applying a standard threshold chosen in the interval $(0,1)$; above the threshold, the correlations are set equal to $1$, below the corresponding opposite value, the correlations are set equal to $-1$, and are 0 in the remaining cases. The binary network loses the values of the edge weights, making the behavior of the correlations sharper and more extreme. The chosen threshold is $0.25$.
The weighted and the binary signed correlation networks are constructed over $\Delta T$-days windows, rolling at $\Delta t$-days step. Then the returns are split by using $\Delta t$-days stepped windows of $\Delta T$-days width and the data from each window are used to construct the correlation matrix of that window. The appropriate choice of $\Delta T$ is a critical point when using eigenvalue-based indicators, since a full rank correlation matrix is required.
@@ 5.1 Comparison between the Global Balance and the Market Rank Indicator
The first database contains the daily returns of the stocks of the S&P500 index, from January 4, 2005 to March 18, 2020. The number of assets is $385$, smaller than the nominal one, since we limited our analysis to the stocks with a complete time series on the whole period. The entire dataset spans $T=4109$ days.
We start with a direct comparison between the behavior of the Global Balance and that of some systemic risk indicators. The time windows width is $\Delta T=400$ days and a step of $\Delta t=30$ days is applied. We note that $\Delta T$ is larger than the number of assets; this permits to calculate the eigenvalue-based indicators. Fig. 5.1 shows the temporal evolution of the Global Balance $\kappa(G)$ and of the average correlation. In Fig 5.1, panels (a) and (b), we have split the entire interval into two sub-intervals, 2005-2016 and 2016-2020, to highlight with higher resolution the information carried by the indicators in the two periods. In Fig. 5.1 panel (c), we zoom in on the period 2008-2014, when the global balance undergoes small relative changes.
The correlation between the two time series is, in general, high. The Pearson correlation is equal to $\rho_{\rm Pearson}=0.5902093$, while rank correlation is equal to $\rho_{\rm Spearman}=0.8771375$.
We also compare the behaviors of the global balance and of the market rank indicator, as shown in Fig. 5.2. Fig. 5.2, panel (a), represents the temporal evolution of the global balance (black line) and of the market rank indicator (blue, red, and green line) over the interval January 2005-October 2016. The three plots for the market rank indicator refer to different choices of the number of smaller eigenvalues used for its computation: the blue line refers to 10 minimum eigenvalues, the red line to 50 and the green line to 100. Similarly, in panel (b), for the period November 2016-January 2020. In the definition of MRI, we have assumed $p=3$ in the power mean used in the denominator (see Eq. (23)). The correlations between the global balance and the MRI for the two intervals and for the different parameters are listed in Table 5.1. Observe that we have added in the Table the case $H=N=385$ that returns the cumulative risk fraction indicator (see [12]).

### v2

@@ 4.3 Approximating Global Balance and Condition Numbers in Correlation Networks
We show now that the ratio ${\mathscr{R}}({\bf A})$, for a large set of significant matrices, in particular for the empirical correlation matrices we are interested in, is less than $1$. We assume from now on that the matrix ${\bf A}$ is positive semidefinite so that all its eigenvalues $\lambda_{i}$ are nonnegative: $0\leq{\lambda}_{N}\leq{\lambda}_{N-1}\leq\dots\leq{\lambda}_{2}\leq{\lambda}_{1}$. We refer to ${\lambda}_{1}$ as the spectral radius and to ${\lambda}_{1}-{\lambda}_{2}$ as the spectral gap.
We first express ${\mathscr{R}}({\bf A})$ as a function of the intervals between pairs of eigenvalues. The ratio ${\mathscr{R}}({\bf A})$ between the condition numbers of the matrices $e^{-{\bf A}}$ and $e^{-|{\bf A}|}$, in the Schatten $p-$norm with $p=1$, can be expressed as
@@
where again ${\lambda}_{i}$ and $\overline{\lambda}_{i},\ i=1,\dots,N$, are the eigenvalues of $\bf A$ and $\bf|A|$, and $\cosh(\cdot)$ denotes the hyperbolic cosine function. The statement follows directly from the identity $\left(\sum_{i=1}^{N}e^{-\lambda_{i}}\right)\left(\sum_{i=1}^{N}e^{\lambda_{i}}\right)=N+2\sum_{j<i}\cosh(\lambda_{i}-\lambda_{j})$.
Formula 19 emphasizes the dependence of ${\mathscr{R}}({\bf A})$ on the intervals $\Delta\lambda_{ij}=\lambda_{i}-\lambda_{j}$ between pairs of eigenvalues. It highlights the relevant role of the spectral gap in large correlation networks based on the correlation matrix ${\bf C}$.
@@
${\mathscr{R}}({\bf A})\approx\frac{1+\frac{2}{N}\sum_{j=2}^{N}\cosh\Delta\lambda_{1j}}{1+\frac{2}{N}\sum_{j=2}^{N}\cosh\Delta\overline{\lambda}_{1j}}\approx\frac{1+\frac{2(N-1)}{N}\cosh\lambda_{1}}{1+\frac{2(N-1)}{N}\cosh\overline{\lambda}_{1}}\approx\frac{e^{\lambda_{1}}}{e^{\overline{\lambda}_{1}}}$ (20)
In general, for any signed network, the inequality $\lambda_{1}\leq\overline{\lambda}_{1}$ holds. The spectral radius of the matrix ${\bf A}$ is strictly smaller than that of the matrix $|{\bf A}|$ if and only if the signed network $G$ is strictly unbalanced, that is neither balanced nor antibalanced, i.e. ${\lambda}_{1}<\overline{\lambda}_{1}$ if and only if $G$ is strictly unbalanced (see [46], Theorem 3.9, and [50]). This implies that the approximated value in Eq. (20) is less than $1$ for strictly unbalanced networks.
Let us observe that, since $\sum_{j=2}^{N}\lambda_{j}=N-\lambda_{1}$ and $\sum_{j=2}^{N}\overline{\lambda}_{j}=N-\overline{\lambda}_{1}$, we have $\sum_{j=2}^{N}\Delta\lambda_{1j}=N(\lambda_{1}-1)<N(\overline{\lambda}_{1}-1)=\sum_{j=2}^{N}\Delta\overline{\lambda}_{1j}$, but this does not guarantees that $\Delta\lambda_{1j}<\Delta\overline{\lambda}_{1j},\ \forall j$. However, if ${\lambda}_{1}\gg{\lambda}_{j}$ and $\overline{\lambda}_{1}\gg\overline{\lambda}_{j}$, for $j=2,\dots,N$, we can also assume $\Delta{\lambda}_{1j}\approx{\lambda}_{1}$ and $\Delta\overline{\lambda}_{1j}\approx\overline{\lambda}_{1}$. This justifies the second approximation in Eq. (20). As a further consequence of the same equation, we can conclude that for correlation networks with large spectral gap the value of the ratio ${\mathscr{R}}({\bf A})$ can be approximated with the value of the global balance $\kappa(G)$, that is ${\mathscr{R}}({\bf A})\approx\kappa(G)$. This effectiveness of the approximation is confirmed by the results in the empirical analysis.
@@
Our interpretation of this result is that the presence of negative edges, that is negative correlations, mitigates the propagation of a perturbation shock within the network. Therefore, an unbalanced graph can mitigate and absorb the perturbations from one node better than a balanced graph. In a strictly unbalanced graph, we expect that the oscillations produced by a single node cannot be amplified to a large extent due to the lower condition number of the adjacency matrix. Conversely, in a balanced graph, where there are two disjoint communities and all closed walks are positive, a perturbation originating from one node is amplified traveling through the network. This result shows that a balanced graph is more unstable and sensitive to perturbations than an unbalanced one, and supports the interpretation of network global balance in terms of predictability in transmitting information, as discussed in subsection 4.1. We conclude this section with a numerical example.
@@
$A=\left[\begin{array}[]{cccc}1&0.764&0.335&0.236\\ 0.764&1&0.051&-0.349\\ 0.335&0.051&1&0.735\\ 0.236&-0.349&0.735&1\\ \end{array}\right].$ (21)
The network associated with this correlation matrix is a weighted version of the graph $K_{4}$ (d) in Fig. 4.1, third line. It is a strictly unbalanced network and its global balance is $\kappa(G)=0.962$. The local balances of the four nodes are collected in the vector $[0.960,0.952,0.983,0.953]$.
Let us first note that by increasing the initial state ${\bf x}_{0}=[1,1,1,1]^{T}$, for example, by the error $\varepsilon{\bf e}_{1}$ with $\varepsilon=2$, the increase in the asymptotic state of node $1$, that is ${\bf x}_{\infty,1}^{\star}-{\bf x}_{\infty,1}=\varepsilon[e^{\bf A}]_{11}$, is equal to $7.641$ for the signed network and to $7.963$ for the unsigned network. The difference for a unitary error, equal to $[e^{|{\bf A}|}]_{11}-[e^{\bf A}]_{11}=0.161$, increases linearly with $\varepsilon$ and the ratio $[e^{\bf A}]_{11}/[e^{|{\bf A}|}]_{11}=0.960$ is exactly the local balance of node $1$. The eigenvalues of ${\bf A}$ are $\lambda_{1}=1.956$ $\lambda_{2}=1.705$. $\lambda_{3}=0.317$ and $\lambda_{4}=0.022$, while the eigenvalues of $|{\bf A}|$ are $\overline{\lambda}_{1}=2.239$ $\overline{\lambda}_{2}=1.274$. $\overline{\lambda}_{3}=0.439$ and $\overline{\lambda}_{4}=0.047$. The two exact condition numbers are $\mathscr{K}(e^{-{\bf A}})=30.378$ and $\mathscr{K}(e^{-{{|\bf A}|}})=30.887$, and their exact ratio is therefore ${\mathscr{R}}({\bf A})=0.984$. For such a low-dimensional correlation matrix, the difference between the two condition numbers is small. Nevertheless it evidences the fact that the balanced network shows a greater dependence on the initial data than the unbalanced network.
This numerical example can be extended to a statistical analysis of the relationship between the global balance $\kappa(G)$ and the ratio of the two condition numbers ${\mathscr{R}}({\bf A})$ in the case of randomly generated correlation matrices. We refer to the Supplementary Material, Sec. 3, for further details.
@@ 4.4 Interpretation in terms of systemic risk measure
Eq. (22) shows that $\lambda_{i}$ represents the portion of the total variance associated with the $i-{th}$ principal portfolio ${\bf\phi}_{i}$, derived from PCA analysis, as defined in [51]. Each component $\lambda_{i}{\bf\phi}_{i}{\bf\phi}_{i}^{T}$ represents a new portfolio out of the original stocks, which is orthogonal to the others. Then the corresponding eigenvalue measures its contribution to the total risk.
@@
where $1\leq H\leq N$ and $p\in{\mathbb{N}}$. Note that the AMRI reduces to the condition number in norm $2$, $\lambda_{1}/\lambda_{N}$, for $H=1$. This measure has been shown to belong to the broad class of so-called proper measures of connectedness ([53]) and have been used to analyze financial datasets in [54]. [^3] Let us observe that, in graphs theory, the quantity $\frac{1}{N}{\rm Tr}\,\bf{C}^{p}=\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}^{p}$ represents the $p-$moment of the adjacency matrix $\bf C$ and, in a positive unweighted network, represents the mean of the number of $p-$cycles around the nodes, that is the average number of $p-$cycles to which a node in the network belongs.
The local and global balance indices introduced in Eqs. (8) and (9) are defined in a way completely independent of both the class of systemic risk measures based on the larger eigenvalues and the class based on the smaller eigenvalues, since they imply a comparison between the correlation matrix and its entry-wise absolute value. However, the interpretation proposed in the previous sections suggests that a bridge can be established between the domain of balance in signed networks and the domain of systemic risk measures for asset correlation networks. Indeed, the relationship of the balance index with the condition numbers of the diffusion problem and with the ratio ${\mathscr{R}}({\bf A})$ defined in Eq. (15) indicates that they are candidates to be indicators of systemic risk as a relative measure of the effect of correlation signs on the spread of information in the underlying network. [^4] It is also possible to show that the global balance trivially satisfies the four properties required in [53] for a proper measure of connectedness. As a consequence, they can be used to detect the greater or lesser instability/sensitivity of the portfolio behavior with respect to the maximum instability/sensitivity associated with the balanced one, which typically has a higher condition number and, consequently, is less diversified and more exposed to systemic risk. At this point, we want to explore the ability of this new indicator to reveal systemic events on empirical correlation networks and compare its capabilities with that of other known measures in the literature.
@@ 5 Empirical analysis
We test the results in Sections 3 and 4 on a real database of financial returns. We focus the analysis on the weighted version of the correlation-based signed network. However, we compare the results with the binary version of the same network. The binarization is obtained by applying a standard threshold chosen in the interval $(0,1)$; above the threshold, the correlations are set equal to $1$, below the corresponding opposite value, the correlations are set equal to $-1$, and are 0 in the remaining cases. The binary network loses the values of the edge weights, making the behavior of the correlations sharper and more extreme. The chosen threshold is $0.25$addedThe choice of this symmetric threshold, although simplifying and inherently arbitrary, is inspired by what was done in Harary’s seminal work [38] and allows a first comparison of the results obtained on the real weighted network and its binarization.
The weighted and the binary signed correlation networks are constructed over $\Delta T$-days windows, rolling at $\Delta t$-days step, that is, the returns are split by using $\Delta t$-days stepped windows of $\Delta T$-days width and the data from each window are used to construct the correlation matrix of that window. The appropriate choice of $\Delta T$ is a critical point when using eigenvalue-based indicators, since a full rank correlation matrix is required.
@@ 5.1 Comparison between the Global Balance and the Market Rank Indicator
The database contains the daily returns of the stocks of the S&P500 index, from January 4, 2005 to March 18, 2020. The number of assets is $385$, smaller than the nominal one, since we limited our analysis to the stocks with a complete time series on the whole period. The entire dataset spans $T=4109$ days. We start with a direct comparison between the behavior of the global balance and that of some systemic risk indicators. The time windows width is $\Delta T=400$ days and a step of $\Delta t=30$ days is applied. We note that $\Delta T$ is larger than the number of assets; this permits to calculate the eigenvalue-based indicators. First, we consider the time series of the global balance and the average correlation in each time window. The correlation between these two time series is generally high. The Pearson correlation is equal to $\rho_{\rm Pearson}=0.5902093$, while rank correlation is equal to $\rho_{\rm Spearman}=0.8771375$. In the Supplementary Material, Sec. 4, we show the time series plots of the global balance and of the average correlation over the period 2005-2020. We now focus on the comparison between the behavior of the global balance and that of the market rank indicator. Their evolution in time are shown in Fig. 5.1.
Fig. 5.1, panel (a), represents the temporal evolution of the global balance (black line) and of the market rank indicator (blue, red, and green line) over the interval January 2005-October 2016. The three plots for the market rank indicator refer to different choices of the number of smaller eigenvalues used for its computation: the blue line refers to 10 minimum eigenvalues, the red line to 50 and the green line to 100. Similarly, in panel (b), for the period November 2016-January 2020. In the definition of MRI, we have assumed $p=3$ in the power mean used in the denominator (see Eq. (23)). The correlations between the global balance and the MRI for the two intervals and for the different parameters are listed in Table 5.1. The last row of the Table ($H=N=385$) returns the special case of the cumulative risk fraction indicator (see [12]).